Q3. Suppose newly diagnosed weekly cases of coronavirus in Wuhan, China (population 11 million) in the first 5 weeks of the epidemic were 10,30,100,400,700.

a) Calculate the cumulative incidence per 100,000 population after 5 weeks (2 points)

b) If the 30 cases in the second week were all secondary cases from the 10 in the first week, what is the basic reproductive number (R0) using just those data? (2 points)

c) Go on to calculate R0 across the first 3 weeks if you assume the 100 cases in the third week were all tertiary cases arising from the 30 in the second week (1 point).

d) Why can the above data not be used to calculate prevalence? (2 points)

e) In the COVID-19 ship outbreak a very high incidence occurred. Explain why this may have been the case and why isolation to cabins did not halt the increasing number of cases. (3 points)
Answer #1

a)Cumulative incidence is the measure of disease frequency during a period of time.

Total number of cases over a period of 5 week=10+30+100+400+700=1240

cumulative incidence=1240/(100000-1240)

=1240/98760

=0.0125.ie,1.25 % over 5 weeks period.

b)Basic Reproductive Number(R0) of an infection is the expected no.of cases directly generated by one case in a population where all individuals are susceptible to infection.

Basic reproductive number=30/10=3.

c)R0=100/30=3.4

d)Prevalence is calculated at a specific point of time wheras incidence over a period of time.As per question,a period of interest is mentioned,not a specific point of time.

e)Control of disease decreases wuth long delays from symptoms onset to isolation,fewer cases ascertained by contact tracing.and increasing transmission before symptoms.

Another important thing is the hesitant natives to come under quarantine,who hides their travel history and health condition.

Before starting strict isolations,contact transmission was happened massively,giving secondary and tertiary cases.These secondary and tertiary cases may not be traced and isolated and thus they remain transmitting;resulting in massive outbreak.